### v3

@@ IV.2 AdS asymptotic region
The firstly one needs to address how the regular solution in Eq.(28) evolves as the asymptotic region is approached, roughly as $r\rightarrow\infty$. As expected, the regular solution, as $r\rightarrow\infty$,
$\lim_{r\rightarrow\infty}f(r)^{2}\approx\gamma+\frac{r^{2}}{l^{2}}-\left(\frac{2M}{r^{d-2n-1}}\right)^{\frac{1}{n}}$ (31)
defining an asymptotically locally AdS space as $r$ grows but also a merging into the vacuum solution counterpart Crisostomo et al. (2000); Aros et al. (2001).
IV.3 Mass
Due to the AdS asymptotia of this family of solutions, the determination of the mass or energy is not straightforward, requiring a regularization process, see for instance Ashtekar and Das (2000); Aros et al. (2001); Skenderis (2002). Here, we use the methodology provided in reference Aros et al. (2001). In this case, given the simple static structure of the solution, any of those methods gives the same result
$E=M+E_{0}$ (32)
where $E_{0}$ vanishes in even dimensions and in odd dimensions stands for the vacuum energy of the AdS space Aros and Estrada (2019).
IV.4 Merging zone
One remarkable feature of the regular solution, see Eq.(28), is the existence, generically, of a value of the radial coordinate where the regular solution converges into the vacuum LnFDGS AdS solution found in references Crisostomo et al. (2000); Aros et al. (2001), which corresponds to Eq.(8).
Analytically, we note that the second factor inside the parentheses in equation (28) is such that:
$\lim_{r/a\rightarrow\infty}\left(1-\exp{\left(-\left(\frac{r}{a}\right)^{\frac{d-1}{n}}\right)}\sum^{n-1}_{i=0}\frac{1}{i!}\left(\frac{r}{a}\right)^{\frac{i}{n}(d-1)}\right)\approx 1,$ (33)
Which implies that for values such that $a\ll r\ll l$, our regular AdS solution and its vacuum AdS counterpart become almost indistinguishable.
However, analytically, it is difficult to estimate from which values of the radial coordinate our regular AdS solution and its vacuum AdS counterpart become indistinguishable. Numerically, we observe in Figure 1 that both cases can be indistinguishable for a finite value of $r_{-}<r=r_{*}<r_{+}$, greater than the inner horizon and less than the event horizon, or equivalently for values of $r$ where $f(r)^{2}<0$. Below, we will discuss some physical implications associated with this fact. It is worth emphasizing that $f(r)^{2}$ can have two, one, or zero roots depending on the values of $d,n,M,a$, and $l$. This is depicted generically in Fig.1.
Figure 1: This is how the merging occurs for different values of the mass parameter. Plot with $r$ on the horizontal axis and $f(r)^{2}$ on the vertical axis. The red curve corresponds to our regular solution, while the blue curve corresponds to the vacuum LnFDGS AdS solutions.
Even though this goes beyond the scope of this work, this rises up the obvious and general question of to what extend a regular solution, black hole or not, can be discerned from its vacuum black hole counterparts by an experiment carried out by an observer beyond that $r$ merging.
IV.5 Structure of horizons
To begin with the discussion it is worthwhile to stress that, and unlike vacuum LnFDGS AdS solutions in Crisostomo et al. (2000); Aros et al. (2001), due to the lack of singularities regular solutions does not need to be black holes, i.e. having a horizon, to be well defined. In other words, regular solutions are not affected by the cosmic censorship hypothesis.
Now rest to answer when a horizon is present for the static regular solution defined by (28). This occurs when $f(r)^{2}$ vanish, i.e., when the set of parameters $a,M$ and $l$ are such that (28) could vanish a given radius.
In figure 2, we can observe the generic behavior of the mass parameter versus radial values where the function $f^{2}(r)$ vanishes. The behavior in the vacuum LnFDGS AdS case Crisostomo et al. (2000); Aros et al. (2001) is illustrated in red, while the behavior of our RBH LUV AdS with our matter sources is shown in blue. Thus, the regions where $M$ is decreasing correspond to the values of the inner horizon, and the regions where $M$ is increasing correspond to the values of the black hole horizon in our solution. Therefore, for $M>M_{min}$, there exists an inner horizon $r_{-}$ and a black hole horizon $r_{+}$ for the same value of the parameter $M$. The value $M_{min}$ corresponds to where the inner and black hole horizons coincide, i.e., an extremal black hole. A notable feature is that for a small deviation to the right in the value of the extremal radius $r_{ext}=r_{-}=r_{+}$ , i.e $r_{*}>r_{ext}$ (which coincides with a small upper deviation of the value of $M$ with respect to $M_{min}$), our regular AdS solution and its vacuum AdS counterpart become indistinguishable, as the red and blue curves are also indistinguishable.
[…]
As a consequence of the above, we also note that for values considerably larger than the Planck scales, such as the observed values of mass and radius of rotating black holes observed in LIGO GW150914 Abbott et al. (2017), both the vacuum and regular solutions are also indistinguishable. Thus, it becomes intriguing to observe the properties of black holes at scales much smaller than those that have been observed to date.
Figure 2: For single value of $M$ there are an interior and exterior horizons. Furthermore, one notice that large exterior $r_{+}$ the vacuum solution is approached
@@ V Shadows
We study the line element in (7) for the spherically symmetric case with $\gamma=1$. Thus, the transversal section can be written as:
@@
$00=\frac{dV_{eff}}{dr}\bigg{|}_{r=r_{sp}}=r_{sp}^{-3}\left(r_{sp}f^{\prime}(r)|_{r=r_{sp}}-2f(r_{sp})\right)\Rightarrow r_{sp}f^{\prime}(r)|_{r=r_{sp}}-2f(r_{sp})=$ 0 (43)
@@
Figure 3: Scheme
Given the lack of an analytical expression for the event horizon, the photon sphere radius, and the shadow size in our case study, and since our case is not asymptotically flat but instead possesses an AdS asymptote, it becomes interesting to establish numerical relationships between the aforementioned shadow characteristics for our case study of Lovelock theories. In the subsection below, we propose a method to achieve this. We will use the case $n=2,d=7$ as an example, but it is straightforward to check that our proposal is applicable to other values of $n$ and $d$.
@@
    As an example, we will study the case $n=2,d=7$ in figure 4. The values of the ascending curve (In the example, as we can see in the orange dashed vertical line, values greater than ∼$\sim 6.5$ on the horizontal axis) represent the candidate values of the photon sphere radius. We use the word ”candidate” because later we need to ensure that the condition $r_{sp}>r_{+}$ is satisfied. The vertical axis shows the values of the mass parameter. Thus, each ordered pair $\left(r_{sp},M(r_{sp})\right)$ represents values of the candidate photon sphere and the mass parameter in the parameter space, such that equation (43) is satisfied, i.e., where the potential reaches an unstable maximum. On the other hand, the orange dashed horizontal line in figure 4 represents the maximum value of the mass parameter for the parameters used. Below, we will detail what the black horizontal line in the figure represents.
    Figure 4: In the horizontal axis the candidate values of the photon sphere radius $r_{sp}$. In the vertical axis the mass parameter $M(r_{sp})$. Using $n=2,d=7,a=4$
@@
    Since equation (1) is independent of the parameter $l$, we can numerically test the values of this parameter where the condition that the shadow radius is greater than the photon sphere radius is satisfied. To do this, we construct the following function in the parameter space:
    $g(r_{sp},l)=\begin{cases}1&\text{if }r_{sh}>r_{sp},\\ 0&\text{if }r_{sh}<r_{sp}.\end{cases}$ (47)
    In this way, we observe in figure 5 that, for the parameters used, the condition $r_{sch}>r_{sp}$ is satisfied for approximately $l>41$ for all value of $r_{sp}$. Therefore, in the next step, we must use a value of $l$ that satisfies the aforementioned condition.
    Figure 5: $g(r_{ph},l)$ using $n=2,d=7,a=4$
@@
    For simplicity, we will write the mass parameter $M(r_{+})$, which depends on the event horizon $r_{+}$, in the parameter space as:
    $M(r_{+})=\frac{r_{+}^{-1+d-2n}\left(1+\frac{r_{+}^{2}}{l^{2}}\right)^{n}}{2\left((-1+n)!-\Gamma\left[n,\left(\frac{r_{+}}{a}\right)^{\frac{d-1}{n}}\right]\right)}$ (48)
    In figure 6, we display the behavior of the mass parameter. We must be consistent with the previous point and choose a value of $l$ that satisfies what was described there. In this way, on the ascending curve, we have an ordered pair of the form $\left(r_{+},M(r_{+})\right)$. The black dashed horizontal line represents the extremal minimum value of the mass for the chosen parameters, emphasizing that $l$ is not arbitrary.
    Figure 6: The values on the horizontal axis of the ascending curve represent the event horizon radius $r_{+}$. On the vertical axis, we have the mass parameter $M(r_{+})$. Using $n=2,d=7,a=4,l=100$.
@@
      The mass parameter must be the same in both cases when the remaining parameters are fixed. Thus, to numerically test the relationship between $r_{+}$ and $r_{sp}$, we must consider that
@@
      Following this condition, the minimum value of the mass parameter is given by the extremal value, represented by the black dashed horizontal line in Figure 6, $M(r_{+})_{min}=M(r_{ext})$ (in the example, ∼$\sim 20$). Therefore, the lower bound for the photon sphere radius is given by $r_{sp}$ such that $M(r_{sp})=M(r_{+})_{min}=M(r_{ext})$ (vertical line in Figure 4, where $r_{sp}\sim 9.8$ and $M(r_{sp}\sim 9.8)\sim 20$). On the other hand, in the orange dashed line of Figure 4, it can be seen that the parameter $M(r_{sp})$ has a maximum value. This last value sets the upper bound for the photon sphere and event horizon, such that condition (49) is satisfied ($r_{sp}\sim 36$ and $r_{+}\sim 23.5$ in the examples of Figures 4 and 6, respectively).
      Once the appropriate ranges for $r_{+}$ and $r_{sp}$ have been determined in the previous steps, we plot the relationship between $r_{+}$ and $r_{sp}$ in Figure 7. This has been done numerically, using condition (49) on the ordered pairs $\left(r_{sp},M(r_{sp})\right)$ and $\left(r_{+},M(r_{+})\right)$ from Figures 4 and 6, respectively. As expected, we observe that the value of the photon sphere is greater than that of the event horizon. That is, photons follow a circular motion due to the geometric distortion of spacetime caused by the presence of the black hole.
      Figure 7: The values on the horizontal axis represent the event horizon radius $r_{+}$. The values on the vertical axis represent the sphere photon radius $r_{sp}$. . Using $n=2,d=7,a=4,l=100$.
@@
    Figure 8: The values on the horizontal axis represent the event horizon radius $r_{+}$. The values on the vertical axis represent the sphere photon radius $r_{sp}$. . Using $n=2,d=7,a=4,l=100$.
@@
Figure 9: The values on the horizontal axis represent the event horizon radius $r_{+}$. The values on the vertical axis represent the sphere photon radius $r_{sp}$. In left panel we have used $n=2,a=4,l=100$, $d=6$(red), $d=7$(blue), $d=8$(black). In right panel we have used $n=3,a=4,l=100$, $d=8$(red), $d=9$(blue), $d=10$(black)
@@
It is also important to mention that in all the cases studied numerically, the condition $r_{+}<r_{sp}<r_{sh}$ is satisfied.
These results could eventually be used for a deeper study in a future paper, estimating, for example, the deflection angle or the shape of a linearly uniformly accelerated trajectory, which can be interpreted as the curved spacetime generalization of the hyperbolic Rindler trajectory in flat spacetime. Both topics are beyond the scope of this work.
Figure 10: The values on the horizontal axis represent the sphere photon radius $r_{sp}$. The values on the vertical axis represent the shadow radius $r_{sh}$. In left panel we have used $n=2,a=4,l=100$, $d=6$(red), $d=7$(blue), $d=8$(black). In right panel we have used $n=3,a=4,l=100$, $d=8$(red), $d=9$(blue), $d=10$(black)
V.2 A brief and speculative glimpse into the M87-EHT constraints
Here, we aim to provide a speculative glimpse into the behavior of the shadow sizes obtained in the previous subsection, in relation to the constraints captured by the EHT, concerning the shadow size of the four-dimensional supermassive black hole M87. Within the $1\sigma$ (68%) confidence level, it can be found that the shadow size of the aforementioned black hole lies within the interval Nozari et al. (2024); Cao et al. (2023); Kocherlakota et al. (2021):
$4.31\leq\frac{2}{V_{\text{vac}}(r=r_{\text{sh}})}=\frac{r_{\text{sh}}}{M}\leq 6.08$ (50)
where $V_{\text{vac}}(r)$ corresponds to the absolute value of the gravitational potential of the vacuum case:
$V_{\text{vac}}(r)=\frac{2M}{r}$ (51)
[…]
$V^{(LL)}_{\text{vac}}(r)=\left(\frac{2M}{r^{d-2n-1}}\right)^{d-2n-1}$ (52)
In this way, and in a speculative manner, we write the ratio of constraint (50) as follows:
$4.31\leq\frac{2^{d-2n-1}}{V^{(LL)}_{\text{vac}}(r=r_{\text{sh}})}=\left(\frac{r_{\text{sh}}^{d-2n-1}}{M}\right)^{1/n}\leq 6.08$ (53)
so, it is direct to check that the ratio $\left(\frac{r_{\text{sh}}^{d-2n-1}}{M}\right)^{1/n}$ is dimensionless and for $n=1,d=4$ the constraint (53) reduces to eq (50).
In Figures 11 and 12, we display the behavior of $\left(\frac{r_{\text{sh}}^{d-2n-1}}{M}\right)^{1/n}$ versus the horizontal axis, which corresponds to a fraction of $M$. In Figure 11, we consider the case $n=2$ with $d=6,7,8$, and in Figure 12, the case $n=3$ with $d=8,9,10$. We observe that, depending on the chosen parameters, the behavior of the vacuum gravitational potential $\frac{2^{d-2n-1}}{V^{(LL)}{\text{vac}}(r=r{\text{sh}})}$ is such that the obtained results can be adjusted to those reported by the EHT for M87 (within the context of four dimensions and General Relativity). In this way, we provide a speculative methodology to compare theoretical results from Lovelock gravity with experimental results from the EHT for M87. Particularly, we find that our results match the experimental results from the EHT for M87 in our plots corresponding to $n=2,\ d=7,8$ and $n=3,\ d=10,11$. It appears that, for a fixed value of $n$, the parameters tend to better fit the experimental results as the number of dimensions increases.
Figure 11: In vertical axis $\left(\frac{r_{\text{sh}}^{d-2n-1}}{M}\right)^{1/n}$ for $n=2,a=4,l=100$. First panel: On the horizontal axis $M/10$ for $d=6$. Second panel: On the horizontal axis $M/100$ for $d=7$. Third panel: On the horizontal axis $M/1000$ for $d=8$
Figure 12: In vertical axis $\left(\frac{r_{\text{sh}}^{d-2n-1}}{M}\right)^{1/n}$ for $n=3,a=4,l=100$. First panel: On the horizontal axis $M/100$ for $d=9$. Second panel: On the horizontal axis $M/1000$ for $d=10$. Third panel: On the horizontal axis $M/10000$ for $d=11$
@@ VI Thermodynamics
From now on, only the black hole solutions, the case where $M>M_{min}$, will be studied. To study the thermodynamics is helpful to express the mass, $M$, in terms of $r_{+}$ as follows
$M(r_{+})=n\left(\gamma+\frac{r_{+}^{2}}{l^{2}}\right)^{n}r_{+}^{d-2n-1}\left(1-\exp{\left(-\left(\frac{r_{+}}{a}\right)^{\frac{d-1}{n}}\right)}\sum^{n-1}_{i=0}\frac{1}{i!}\left(\frac{r_{+}}{a}\right)^{\frac{i}{n}(d-1)}\right)^{-1}$ (54)
Analytically, we note that for $r_{+}/a\rightarrow\infty$, this mass function converges to the vacuum solution. However, as mentioned above, from Figure 2 we can observe numerically that for a value of the event horizon smaller than the one mentioned, which is slightly greater than the extremal radius, $r_{*}>r_{ext}$, the mass parameter of both cases becomes indistinguishable.

### v4

@@ IV.4 AdS asymptotic region
The firstly one needs to address how our regular solution, equations (25) and (26), evolves as the asymptotic region is approached, roughly as $r\rightarrow\infty$. As expected, the regular solution, as $r\rightarrow\infty$,
$\lim_{r\rightarrow\infty}f(r)\approx 1+\frac{r^{2}}{l^{2}}-\left(\frac{2M}{r^{d-2n-1}}\right)^{\frac{1}{n}}$ (31)
defining an asymptotically locally AdS space as $r$ grows but also a merging into the vacuum solution counterpart Crisostomo et al. (2000); Aros et al. (2001). As is well known, an asymptotically AdS behavior allows for a proper definition of conserved quantities at the boundary Aros et al. (2000). Along these lines, shadows have also been studied in spacetimes with this type of asymptotics Belhaj et al. (2021); Hamil et al. (2024).
@@ V Shadows
We study the spherically symmetric line element given by (7). Thus, the transversal section can be written as:
@@
$0=\frac{dV_{eff}}{dr}\bigg|_{r=r_{sp}}=r_{sp}^{-3}\left(r_{sp}f^{\prime}(r)|_{r=r_{sp}}-2f(r_{sp})\right)\Rightarrow r_{sp}f^{\prime}(r)|_{r=r_{sp}}-2f(r_{sp})=0$ (41)
@@
Figure 3: Scheme: We observe that photons launched with an impact parameter smaller than the shadow radius fall into the black hole. Photons with an impact parameter equal to the shadow radius fall into an unstable circular orbit. Photons launched with an impact parameter greater than the shadow radius escape toward the boundary of spacetime.
As previously mentioned, the study of black holes in Lovelock gravity has attracted significant attention in recent years from a theoretical perspective. This naturally motivates the investigation of their shadow properties. In this regard, given the lack of analytical expressions for the event horizon, the photon sphere radius, and the shadow size in our case study, it becomes particularly interesting to establish numerical relationships between these shadow-related features within the context of Lovelock theories. In what follows, we propose a numerical methodology to achieve this. In the subsection below, we propose a method to achieve this. We will use the case $n=2,d=7$ as an example, but it is straightforward to check that our proposal is applicable to other values of $n$ and $d$.
@@
    As an example, we will study the case $n=2,d=7$ in figure 4. The values of the ascending curve (In the example, as we can see in the orange dashed vertical line, values greater than $\sim 6.5$ on the horizontal axis) represent the candidate values of the photon sphere radius. We use the word ”candidate” because later we need to ensure that the condition $r_{sp}>r_{+}$ is satisfied. The vertical axis shows the values of the mass parameter. Thus, each ordered pair $\left(r_{sp},M(r_{sp})\right)$ represents values of the candidate photon sphere and the mass parameter in the parameter space, such that equation (41) is satisfied, i.e., where the potential reaches an unstable maximum. On the other hand, the orange dashed horizontal line in figure 4 represents the maximum value of the mass parameter for the parameters used. Below, we will detail what the black horizontal line in the figure represents.
    Figure 4: This figure displays the range of possible values for the photon sphere. In the horizontal axis the candidate values of the photon sphere radius $r_{sp}$. In the vertical axis the mass parameter $M(r_{sp})$. Using $n=2,d=7,a=4$
@@
    We are interested in testing the parameter values that satisfy the condition that the shadow radius is greater than the photon sphere radius, $r_{sh}>r_{sp}$. Due to the highly nonlinear nature of our equations, obtaining analytical relations for this condition becomes extremely difficult.
    From equation (1) and figure 4, we extract the ordered pairs $\left(r_{sp},M(r_{sp})\right)$, which allow us to identify the possible values of the photon sphere radius. On the other hand, from equation (44), we obtain the corresponding shadow radius values. Using this information, we determine the potential values of both the photon sphere radius and the shadow radius as given by equation (44).
    However, as previously mentioned, we are interested in identifying the values of the parameter $l$ that satisfy the condition that the shadow radius is greater than the photon sphere radius. To this end, using equations (1) and (44), we employ the RegionPlot command in Mathematica. This is facilitated by the fact that equation (1) is independent of the parameter $l$.
    As shown in figure 5, for the parameters considered, the condition $r_{sch}>r_{sp}$ is fulfilled for approximately $l>41$ across all values of $r_{sp}$. Therefore, in the following step, we must use a value of $l$ that satisfies this condition.
    Figure 5: In blue, we observe the values of the photon sphere radius $r_{\text{sp}}$ and the AdS radius $l$ that satisfy the condition $r_{\text{sh}}>r_{\text{sp}}$
    , using $n=2,d=7,a=4$
@@
    To display the numerical behavior of the mass parameter, we use Equation (28), where, as previously mentioned, the values on the horizontal axis in the ascending curve correspond to the values taken by the event horizon, $\bar{r}=r_{+}$.
    In figure 6, we display the behavior of the mass parameter. We must be consistent with the previous point and choose a value of $l$ that satisfies the description provided there. In this way, on the ascending curve, we have an ordered pair of the form $\left(r_{+},M(r_{+})\right)$. The black dashed horizontal line represents the extremal minimum value of the mass for the chosen parameters, emphasizing that $l$ is not arbitrary.
    Figure 6: This figure displays the values of the event horizon corresponding to different values of the mass parameter. The values on the horizontal axis of the ascending curve represent the event horizon radius $r_{+}$. On the vertical axis, we have the mass parameter $M(r_{+})$. Using $n=2,d=7,a=4,l=100$.
@@
    • The mass parameter must be the same in both cases when the remaining parameters are fixed. This is because $M$ represents the mass parameter of the solution and must have the same value for both $r_{sp}$ and $r_{+}$. Thus, to numerically test the relationship between $r_{+}$ and $r_{sp}$, we must consider that
@@
    • Following this condition, the minimum value of the mass parameter is given by the extremal value, represented by the black dashed horizontal line in Figure 6, $M(r_{+})_{min}=M(r_{ext})$ (in the example, $\sim 20$). Therefore, the lower bound for the photon sphere radius is given by $r_{sp}$ such that $M(r_{sp})=M(r_{+})_{min}=M(r_{ext})$ (vertical line in Figure 4, where $r_{sp}\sim 9.8$ and $M(r_{sp}\sim 9.8)\sim 20$).
      On the other hand, as shown by the orange dashed line in Figure 4, the parameter $M(r_{sp})$ reaches a maximum value. This last value sets the upper bound for the photon sphere and event horizon, such that condition (45) is satisfied ($r_{sp}\sim 36$ and $r_{+}\sim 23.5$ in the examples of Figures 4 and 6, respectively).
    • Once the appropriate ranges for $r_{+}$ and $r_{sp}$ have been determined in the previous steps, we plot the relationship between $r_{+}$ and $r_{sp}$ in Figure 7. This has been done numerically, using condition (45) on the ordered pairs $\left(r_{sp},M(r_{sp})\right)$ and $\left(r_{+},M(r_{+})\right)$ from Figures 4 and 6, respectively. As expected, we observe that the value of the photon sphere is greater than that of the event horizon. That is, photons follow a circular motion due to the geometric distortion of spacetime caused by the presence of the black hole. We also note that as $r_{+}$ increases, the value of $r_{sp}$ also increases in the parameter space.
      Figure 7: This figure displays the values of the photon sphere radius versus the values of the event horizon radius. The values on the horizontal axis represent the event horizon radius $r_{+}$. The values on the vertical axis represent the sphere photon radius $r_{sp}$. . Using $n=2,d=7,a=4,l=100$.
@@
    Figure 8: This figure displays the values of the shadow radius corresponding to different values of the photon sphere radius. Using $n=2,d=7,a=4,l=100$.
@@
Figure 9: This figure displays the values of the photon sphere radius corresponding to different values of the event horizon radius. The values on the horizontal axis represent the event horizon radius $r_{+}$. The values on the vertical axis represent the sphere photon radius $r_{sp}$. In left panel we have used $n=2,a=4,l=100$, d=6d=6(red), d=7d=7(blue), d=8d=8(black). In right panel we have used $n=3,a=4,l=100$, d=8d=8(red), d=9d=9(blue), d=10d=10(black)
@@
From a physical point of view, our methodology allows us to estimate the parameter ranges for which the condition that the shadow radius fully encloses the photon sphere radius, and the latter, in turn, fully encloses the event horizon, is satisfied , that is, the condition $r_{+}<r_{\text{sp}}<r_{\text{sh}}$ holds. This is consistent with the analytical treatment of the vacuum Pure Lovelock case presented in Reference Paithankar and Kolekar (2023), which appears to be the only analytically solvable case currently known for Lovelock gravity.
In this way, our numerical approach could eventually be used for a deeper investigation in a future work, enabling, for example, the estimation of the deflection angle or the characterization of a linearly uniformly accelerated trajectory. In addition to the above, we would like to point out that establishing a direct comparison between the AdS radius values used in our figures and those estimated by other authors in the context of Lovelock theories would require , at the very least, theoretical justification regarding the behavior of the full set of black hole parameters within the AdS–Lovelock framework. This latter aspect goes beyond the scope of the present work. As mentioned, the values of the AdS radius currently available in the literature for Lovelock gravity remain speculative Chernicoff et al. (2016); Camanho et al. (2016).
Figure 10: This figure displays the values of the shadow radius corresponding to different values of the photon sphere radius. The values on the horizontal axis represent the sphere photon radius $r_{sp}$. The values on the vertical axis represent the shadow radius $r_{sh}$. In left panel we have used $n=2,a=4,l=100$, d=6d=6(red), d=7d=7(blue), d=8d=8(black). In right panel we have used $n=3,a=4,l=100$, d=8d=8(red), d=9d=9(blue), d=10d=10(black)
@@ VI Thermodynamics
First of all, it is worth noting that the energy can be computed as a conserved charge by following the methodology in Reference Aros et al. (2001), which involves a regularization process. In this case, the conserved charge associated with the energy, computed using a timelike Killing vector, is obtained in an analogous way to Reference Aros and Estrada (2019):
$E=M$ (46)
In this way, the black hole energy is directly related to the parameter $M$.
On the other hand, in Reference Estrada (2022) it was shown, using the conditions $f(r_{+},M)=0$ and $\delta f(r_{+},M)=0$, which can be viewed as constraints on the evolution along the space of parameters, that the first law of thermodynamics takes the form:
$\frac{\partial{m}}{\partial M}dM=\left(\frac{1}{4\pi}f^{\prime}\big|_{r=r_{+}}\right)\left(2\pi nr_{+}^{d-2n-1}\left(1+\frac{r_{+}^{2}}{l^{2}}\right)^{n-1}dr_{+}\right).$ (47)
The above equation can be rewritten as:
$du=TdS,$ (48)
where we can identify the temperature and entropy terms as:
$T=$ $\frac{1}{4\pi}f^{\prime}\big|_{r=r_{+}}$ (49)
$dS=$ $2\pi nr_{+}^{d-2n-1}\left(1+\frac{r_{+}^{2}}{l^{2}}\right)^{n-1}dr_{+}$ (50)
where $du$ corresponds to a local definition of the variation of the energy at the horizon. On the other hand, the definitions of entropy and temperature coincide with those of the vacuum case. That is, our matter sources influence the behavior of the aforementioned definition $du$.
In this section, we study the thermodynamics at the event horizon, that is, where the mass parameter satisfies $M\geq M_{\text{min}}$, where $M_{\text{min}}$, as previously mentioned, is associated with the extremal radius, where the inner and event horizons coincide, $r_{\text{ext}}=r_{-}=r_{+}$. It is worth emphasizing, as mentioned earlier, that an interesting feature of our solution is that, for a value of $M$ slightly greater than $M_{\text{min}}$, the behavior of the mass parameter in our solution becomes indistinguishable from the vacuum case. Therefore, under this condition, the event horizons of both solutions coincide. As also discussed, this implies that the geometry of both cases becomes indistinguishable beyond the event horizon.